A river, 400m wide is flowing with a rate of 2 m/s. A boat is sailing at a velocity of 10m/s with respect to water, in a direction perpendicular to the river. The time taken by boat to reach the opposite bank is?
1
Expert's answer
2018-11-19T16:00:09-0500

Since a boat is sailing in a direction perpendicular to the river, the time needed to reach the opposite bank does not depend on velocity of river flow. So

t=L/v=400/10=40s